# Scalable Machine Learning Models for Predicting Quantum Transport in Disordered 2D Hexagonal Materials

### v1

@@ 3 Results
Similarly, Table 3 shows that regression is also superior for $LDOS$ predictions, with MAE=0.0085, whereas classification achieves an MAE of 0.1002. The classification model struggles with capturing finer variations in LDOS due to discretization effects. Discretizing continuous values by rounding leads to information loss and this approach affects on performance of RF classification. While classification models achieve high accuracy and F1-score, they suffer from information loss due to rounding. This explains why regression models consistently achieve higher scores. Classification forces continuous values into discrete bins, making it unable to capture small variations in T and $LDOS$ that regression can learn directly.
Figure 5: predict T and LDOS values for random configuration from the test dataset
(fig. 5) shows the results of using RF regression as a selected model in one of the random configurations in the test dataset. The results agree with the performance metrics.
@@
To evaluate the models’ performance on extrapolation data—system configurations outside the training range, we generated a test dataset in the systems shape of: $w:36,L:2,nm:5$ and predict targets with our models.
@@
Figure 6: predict T and LDOS values for extrapolation data of graphene system
Table 4 presents the performance of our models on extrapolation data. Significant decrease in performances are observed: the accuracy of the RF regression model for T drops by $32\%$, and for $LDOS$, it drops by $25\%$. The overall degradation in performance across all metrics highlights a fundamental limitation of RF models. (fig. 6) is the predicted values versus reals of graphene sample where shows that the Random Forest model struggles with extrapolation. While it performs well on training and test data, its accuracy drops significantly when applied to configurations outside the training range. This is because RF builds decision trees based on thresholds observed during training, and when features exceed these thresholds, the model lacks meaningful rules to guide predictions. As a result, it often defaults to predictions from boundary regions of the training set, leading to poor generalization.
Random Forest (RF) has been successfully applied in previous works to classify T, demonstrating its ability to capture transport patterns[15] [38]. However, such models typically focus on small-scale systems or limited feature sets, and do not address scalability across material types and device geometries. In this work, we extend the analysis by testing RF in both regression and classification settings while ensuring the model can generalize to larger system sizes and impurity levels. Importantly, we assess the models’ limitations by testing them on extrapolation data — a critical step often missing in similar works. These results clearly demonstrate both the practical value and boundaries of Random Forest models in quantum transport prediction.
@@
Here, we built data-driven and scalable machine learning models to predict key quantum transport quantities—namely, the transmission coefficient T(E) and the local density of states (LDOS)—in hexagonal two-dimensional (2D) lattices with randomly located magnetic impurities. The model for the physical system was a tight-binding Hamiltonian, and the transport quantities were computed using the Non-Equilibrium Green’s Function (NEGF) formalism. By constructing a dataset that spans a broad range of device geometries, impurity concentrations, and energy levels, we ensured that our models learned from physically valid and diverse configurations. Features were selected based on the geometric and physical characteristics of the lattice; including the width and length of the scattering region, number of impurities, hopping amplitude, and target energy. Our results reveal that Random Forest regression models are superior to classification-based approaches in accuracy, stability, and the ability to learn continuous variations in quantum transport behavior. Regression models possessed extremely low error measures and very high $R^{2}$ scores on test data, successfully learning the underlying nonlinear relationships created by impurity-induced disorder and interference effects.
@@
In the future, one can look forward to seeing further work in the direction of merging more expressive and adaptive learning frameworks—such as Physics-informed neural networks (PINNs) , graph neural networks, or Gaussian processes—to improve generalization, especially in extrapolation regimes. Further, opening up the input feature space to include spin-orbit coupling, edge geometries, temperature effects, or even strain and external fields would help to further improve the physical accuracy and predictive power of the model. These developments would allow the models to be used for a broader class of materials and device geometries, ultimately pushing the limits of machine learning in quantum transport.

### v2

@@ 3 Results
Similarly, Table 4 shows that regression is also superior for average-LDOS predictions, with MAE = 0.006, whereas classification achieves an MAE of 0.064. The classification model struggles with capturing finer variations in average-LDOS due to discretizations effects. Discretizing continuous values by rounding leads to information loss and this approach affects the performance of RF classification. Classification forces continuous values into discrete bins, making it unable to capture small variations in $T$ and average-LDOS that regression can learn directly. Additionally, for predicting the average-LDOS, models with polynomially expanded features also exhibited superior performance compared to those without feature expansion.
Figure 5: Regression model performance. Comparing the predicted vs. true values for Average-LDOS and Transmission ($T$). The dashed red line represents the ideal prediction ($y=x$), and the shaded area shows the $95\%$ Prediction Interval (PI).
The comparative performance of the regression model across the two target properties, Transmission and Average-LDOS, is visualized in Fig 5, which plots the predicted values against the true values alongside the $95\%$ Prediction Interval (PI). The predictions for Transmission ($T$) exhibit exceptionally high precision, clustering tightly around the ideal $y=x$ line with a notably narrow $95\%$ PI across the full range of values . This visual agreement confirms the model’s high predictive power for the global transport property.In contrast, the predictions for Average-LDOS show a marked increase in uncertainty and scatter, particularly evident in the wider and more volatile $95\%$ PI . Furthermore, a systematic bias emerges at higher Average-LDOS values (above $\approx 2.0$), where the model consistently underpredicts the true response. his increased prediction error for the local Average-LDOS property highlights its inherent complexity and greater sensitivity to fine geometric variations, emphasizing why approaches like polynomial feature expansion were necessary to achieve even this level of accuracy. Fig.6 complements this analysis by showing the full spectral prediction for both Average-LDOS and Transmission, $T(E)$, as a function of energy ($E$) for one random configuration in the test dataset. The fidelity of the predicted spectral curves (orange dots) relative to the real NEGF calculated values (blue dots) visually confirms the quantitative findings. The model successfully reproduces the intricate features of both $T(E)$ and Average-LDOS spectra, with the $95\%$ PI (green shaded area) accurately encompassing the real values at nearly all energy points.
Figure 6: Predicted (orange) and Real (blue) spectral curves for Average-LDOS and Transmission $T(E)$ as a function of energy ($E$), shown for one representative configuration from the test dataset. The green shaded region indicates the $95\%$ Prediction Interval (PI).
@@
To evaluate the models performance on extrapolation data, i.e., system configurations outside the training range, we generated a test dataset by graphene, stanene, silicene, and germanene; in the system shapes of: $w:14,36$ ; $L:5,9$ ; $nm:6\%$ and predict targets with our models.
@@
Figure 7: Regression model performance on extrapolation dataset. Comparing the predicted vs. true values for Average-LDOS and Transmission ($T$) . The dashed red line represents the ideal prediction ($y=x$), and the shaded area shows the $95\%$ Prediction Interval (PI).
Table 5 presents the performance of our models on extrapolation data. Significant decreases in performance are observed: the accuracy of the RF regression model for $T$ drops by $38\%$, and for Average-LDOS, it drops by $25\%$. The evaluation of the model on the extrapolation dataset, visualized in Fig.7, reveals the inherent limitations of generalization outside the training domain, particularly in the model’s confidence. For Transmission, the model maintains reasonable accuracy along the $y=x$ line, demonstrating robustness for the global transport property even in highly disordered and short systems. However, the $95\%$ Prediction Interval (PI) significantly widens compared to the in-domain test, reflecting a substantial loss in prediction certainty as the model moves away from familiar configurations.Also for Average-LDOS, the extrapolation performance degrades. The true Average-LDOS values in this test are in the low range ($\approx 0.0$ to $1.4$), a region which showed the highest certainty in the in-domain analysis. Nevertheless, PI becoming extremely volatile and widely dispersed across the entire plot and performance decrease due of novel geometric inputs.
The overall degradation in performance across all metrics highlights a fundamental limitation of RF models. Fig.8 is the predicted values versus reals of graphene sample where shows that the Random Forest model struggles with extrapolation. While it performs well on training and test data, its accuracy drops significantly when applied to configurations outside the training range. This is because RF builds decision trees based on thresholds observed during training, and when features exceed these thresholds, the model lacks meaningful rules to guide predictions. As a result, it often defaults to predictions from boundary regions of the training set, leading to poor generalization.
Figure 8: Predict $T$ and Average-LDOS values for extrapolation data of graphene system.
Random Forest (RF) has been successfully applied in previous works to classify $T$, demonstrating its ability to capture transport patterns[Ghosh2023-pc] [9216557]. However, such models typically focus on small-scale systems or limited feature sets, and do not address scalability across material types and device geometries. In this work, we extend the analysis by testing RF in both regression and classification settings while ensuring the model can generalize to larger system sizes and impurity levels. Importantly, we assess the models limitations by testing them on extrapolation data, a critical step often missing in similar works. These results clearly demonstrate both the practical value and boundaries of Random Forest models in quantum transport prediction.
@@
Here, we built data-driven and scalable machine learning models to predict key quantum transport quantities namely, the transmission coefficient $T(E)$ and the local density of states ($LDOS$) in hexagonal two-dimensional (2D) lattices with randomly located magnetic impurities. The model for the physical system was a tight-binding Hamiltonian, and the transport quantities were computed using the Non-Equilibrium Green’s Function (NEGF) formalism. By constructing a dataset that spans a broad range of device geometries, impurity concentrations, and energy levels, we ensured that our models learned from physically valid and diverse configurations. Features were selected based on the geometric and physical characteristics of the lattice; including the width and length of the scattering region, number of impurities, hopping amplitude, and target energy. Our results reveal that Random Forest regression models are superior to classification-based approaches in accuracy, stability, and the ability to learn continuous variations in quantum transport behavior. Regression models possessed extremely low error measures and very high $R^{2}$ scores on test data, successfully learning the underlying nonlinear relationships created by impurity-induced disorder and interference effects.
@@
In the future, one can look forward to seeing further work in the direction of merging more expressive and adaptive learning frameworks such as Physics-informed neural networks (PINNs), graph neural networks, or Gaussian processes to improve generalization, especially in extrapolation regimes. Further, opening up the input feature space to include spin-orbit coupling, edge geometries, temperature effects, or even strain and external fields would help to further improve the physical accuracy and predictive power of the model. These developments would allow the models to be used for a broader class of materials and device geometries, ultimately pushing the limits of machine learning in quantum transport.
Data and Code Availability
The machine learning code developed for this study and the dataset used to train and evaluate the models are publicly available in the Zenodo repository [seyed_mahdi_mastoor_2025_17109064]. The corresponding live repository for code development is available on GitHub
Conflicts of interest
There are no conflicts to declare.
Author Contributions
Seyed Mahdi Mastoor: investigation, validation, software, programming, writing – original draft, writing – reviewing, editing, resources and Amirhossein Ahmadkhan Kordbacheh: supervision, project administration, writing – reviewing, editing, investigation.